# SG-LRA: Self-Generating Automatic Scoliosis Cobb Angle Measurement with Low-Rank Approximation

### v1

@@ 4.1.1 Datasets
We evaluate Cobb angle measurement on public AASCE2019 [11, 60] benchmark, and our private Spinal2023 and open-source Spinal-AI2024 datasets. The dataset details are summarized in Table 1.
@@
  • SpinalAI2024 is our generated dataset with totally 20,000 spine X-ray images, in which 16,000 and 4,000 images are used for training and testing, respectively. This dataset has a broad distribution of Cobb angles with a certain range of image intensities. The annotations of spinal landmarks as well as three Cobb angles for each image is obtained by our proposed data engine.
@@ 4.1.2 Implementation Details
Our SG-LRA is implemented via PyTorch. Similar to previous work [20], we set the eigen-spine space dimension $M$ and the spinal segment vertex number $N$ to 16 and $14$, respectively. The loss weights $\lambda_{reg}$ and $\lambda_{cls}$ in Eq. (5) are set to 0.1 and 1, respectively. We adopt stochastic gradient descent (SGD) optimizer and polynomial learning rate scheduler, with a weight decay of 0.0005, a momentum of 0.9, and an initial learning rate of 0.001. During pseudo-labeling, our spinal contour detection network is first pre-trained for 100 epochs on Spinal2023, and then are iteratively fine-tuned on training data composed of Spinal2023 images and selected generated images. The confidence threshold $\tau_{c}$ is set to 0.3. For privacy review, the weights $\lambda_{ss}$ and $\lambda_{ps}$ in Eq. (13) are set as 0.2 and 0.8, and the CS threshold $\tau_{CS}$ is set as 0.6.
@@ 4.1.3 Evaluation Metrics
Table 2: Comparison with state-of-the-art methods on AASCE2019 in terms of maximum Cobb angle evaluation. These methods are divided into traditional, regression-based, segmentation-based, and landmark-based methods, respectively. Our SG-LRA can be regarded as combining regression-based and landmark-based methods.
@@ 4.2.1 Maximum Cobb Angle Evaluation
Table 2 shows comparison results on the public AASCE2019 dataset. It can be seen that our SG-LRA maintains top-tier performance on AASCE2019. Compared to regression-based methods like Faster-RCNN, SG-LRA performs better, due to our LRA parameterized spine representation for precise Cobb angle calculation. Compared to segmentation-based methods, SG-LRA with sparse and dense assignments can more accurately detect spinal segment contours, better addressing issues of spinal mask connectivity/fragmentation. Compared to landmark-based methods, our LRA coefficient regression combined with the anchor box classification branch can more accurately localize corner points, giving stronger robustness for non-rectangular spinal segments.
@@
Besides, we visualize different spine representations in Fig. 3. It can be seen that our LRA parameterized spine representation is more fine-grained than landmark-based and rectangular-box-based representations, enabling more precise Cobb angle calculation.
@@ 4.2.2 PT, MT, and TL/L Cobb Angle Evaluation
We conduct a more detailed comparison by evaluating Cobb angles of three regional spinal curves in Table 3. Since the construction of Spinal-AI2024 involves our spinal contour detection network, the reported results of SG-LRA on Spinal-AI2024 are obtained by re-training spinal contour detection network on Spinal-AI2024. It can be observed that our SG-LRA significantly outperforms previous works in most cases.
In particular, on all the three datasets, SG-LRA performs the best on MT and TL/L Cobb angles, as well as maximum Cobb angle. This is mainly thanks to the combination of LRA coefficient regression and anchor box classification. Note that the performance for PT cure is not as good, possibly because SG-LRA focuses more on localizing irregular spinal segments rather than limiting the number of predicted spinal segments. As a result, it may involve the cervical spine in the calculation of PT Cobb angle, leading to worse results. Although previous methods fix the predicted whole scoliosis to 17 spinal segments, which often obtains more accurate PT predictions, complex pre-processing of data and annotations based on clinical knowledge are required.
It can be seen that segmentation-based method Residual U-Net achieves the best PT Cobb angle estimation result on Spinal2023, and landmark-based method Linformer achieves the best PT Cobb angle estimation result on Spinal-AI2024. We notice that Spinal2023 has limited images while Spinal-AI2024 has a larger scale. In this case, landmark-based methods suffer from insufficient training data on Spinal2023 but can be trained well on Spinal-AI2024, while segmentation-based methods have larger advantages on Spinal2023 with smaller scale. In contrast, our SG-LRA consistently works well across small and large datasets.
@@ 4.3.1 Spinal Contour Detection Network
Our spinal contour detection network consists of LRA coefficient regression branch and anchor box classification branch, with each branch containing sparse and dense assignments. Table 4 presents the results of different variants on Spinal2023.
@@ 4.3.2 Data Engine
Table 6: Results using different selections on Spinal-AI2024 from the first to the six iterations. “No”, “Indep.”, and “Cumul.” denote no filtering, and updating $V_{selected}$ in independent and cumulative ways across iterations, respectively.
@@
Different Sample Selection Strategies. We conduct three experiments using different sample selection strategies, as shown in Table 6. If a sample is filtered and will not be selected in the next iterations, i.e. cumulative way, the overall highest quality improvement can be achieved. If $V_{selected}$ is updated independently across iterations, the quality improvement becomes worse but is still better than no filtering. This demonstrates the cumulative effectiveness of combined denoising and privacy review processes.
@@ 5.2.3 Generated Spinal-AI2024
For each Spinal-AI2024 image, a top-3 comprehensive similarity review is presented in Table 7, in which the most similar images are shown in Fig. 6. The results show that our generated dataset has distorted the original annotations and produced spinal curvature variations, which can prevent excessive patient information leakage.
@@ 6 Conclusion
We have proposed a novel self-generating framework utilizing LRA parameterized spine representation for automatic Cobb angle measurement. LRA coefficient regression and anchor box classification with sparse and dense assignments are combined to detect irregular contours of spinal segments. Besides, we have proposed a data engine with semi-supervised labeling and privacy review, which culminates in the open-source largest scoliosis X-ray dataset Spinal-AI2024. Extensive experiments have demonstrated the effectiveness of our method, achieving state-of-the-art performance on public, private, and generated datasets for both maximum Cobb angle and three regional Cobb angles.

### v2

@@ 4.1.1 Datasets
We evaluate Cobb angle estimation on public AASCE2019 [5, 39] benchmark, and our private Spinal2023 and open-source Spinal-AI2024 datasets. The dataset details are summarized in Table 1.
  • 1.
@@
  • 3.
    Spinal-AI2024 is our generated dataset with totally 20,000 spine X-ray images, in which 16,000 and 4,000 images are used for training and testing, respectively. This dataset has a broad distribution of Cobb angles with a certain range of image intensities. The annotations of spinal landmarks as well as three Cobb angles for each image is obtained by our proposed data engine.
@@ 4.1.2 Implementation Details
Our CurvNet is implemented via PyTorch on an NVIDIA GeForce RTX 4090 GPU. Similar to previous work [11], we set the eigen-spine space dimension $M$ and the spinal segment vertex number $N$ to 16 and $14$, respectively. The eigen-spines are constructed offline via SVD on Spinal2023 by taking 6.2 seconds, and are reused across all inference stages, ensuring good scalability. The loss weights $\lambda_{reg}$ and $\lambda_{cls}$ in Eq. (5) are set to 0.1 and 1, respectively. We adopt stochastic gradient descent (SGD) optimizer and polynomial learning rate scheduler, with a weight decay of 0.0005, a momentum of 0.9, and an initial learning rate of 0.001. During pseudo-labeling, our spinal contour detection network is first pre-trained for 100 epochs on Spinal2023, and then are iteratively fine-tuned on training data composed of Spinal2023 images and selected generated images. The confidence threshold $\tau_{c}$ is set to 0.3. For privacy review, the weights $\lambda_{ss}$ and $\lambda_{ps}$ in Eq. (13) are set as 0.2 and 0.8, and the CS threshold $\tau_{CS}$ is set as 0.6. During inference, our spinal contour detection network runs at 16.7 frames per second (FPS), suitable for real-time clinical assistance.
For spinal image generation, we fine-tune the U-Net of Stable Diffusion v1.5 for 50 epochs using the AdamW optimizer with $\text{lr}=2\times 10^{-5}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The used text prompts include “AP view of human spine X-ray, scoliosis, clear vertebrae, no artifacts”. We also utilize random horizontal flips and intensity jitter for data augmentation. Latent codes are sampled from $\mathcal{N}(0,I)$, and classifier-free guidance with scale=7.5 is used to balance realism and diversity.
@@ 4.1.3 Evaluation Metrics
Table 2: Comparison with state-of-the-art methods on AASCE2019 in terms of maximum Cobb angle evaluation. These methods are divided into traditional, regression-based, segmentation-based, and landmark-based methods, respectively. Our CurvNet can be regarded as combining regression-based and landmark-based methods.
@@ 4.2.1 Maximum Cobb Angle Evaluation
Table 2 shows comparison results on the public AASCE2019 dataset. It can be seen that our CurvNetmaintains top-tier performance on AASCE2019. Compared to regression-based methods like Faster-RCNN, CurvNet performs better, due to our latent contour parameterized spine representation for precise Cobb angle calculation. Compared to segmentation-based methods, CurvNet with sparse and dense assignments can more accurately detect spinal segment contours, better addressing issues of spinal mask connectivity and fragmentation. Compared to landmark-based methods, our latent contour coefficient regression combined with the anchor box classification branch can more accurately localize corner points, giving stronger robustness for non-rectangular spinal segments.
@@
Besides, we visualize different spine representations in Fig. 3. It can be seen that our latent contour parameterized spine representation is more fine-grained than landmark-based and rectangular-box-based representations, enabling more precise Cobb angle calculation.
@@ 4.2.2 PT, MT, and TL/L Cobb Angle Evaluation
We conduct a more detailed comparison by evaluating Cobb angles of three regional spinal curves in Table 3. Since the construction of Spinal-AI2024 involves our spinal contour detection network, the reported results of CurvNet on Spinal-AI2024 are obtained by re-training spinal contour detection network on Spinal-AI2024. It can be observed that our CurvNet significantly outperforms previous works in most cases.
In particular, on all the three datasets, CurvNet performs the best on MT and TL/L Cobb angles, as well as maximum Cobb angle. This is mainly thanks to the combination of latent contour coefficient regression and anchor box classification. Note that the performance for PT cure is not as good, possibly because CurvNet focuses more on localizing irregular spinal segments rather than limiting the number of predicted spinal segments. As a result, it may involve the cervical spine in the calculation of PT Cobb angle, leading to worse results. Although previous methods fix the predicted whole scoliosis to 17 spinal segments, which often obtains more accurate PT predictions, complex pre-processing of data and annotations based on clinical knowledge are required.
To more thoroughly investigate the failure causes of PT Cobb angle estimation, we analyze the failure cases by our CurvNet on the test set of AASCE2019. We define a PT estimation error as “large” if the absolute difference between predicted and ground-truth angle exceeds $10^{\circ}$. There are 37 of 128 test images with large estimation errors:
  • 1.
    29 cases (78%) involve the erroneous detection of cervical vertebrae (C1–C7) as part of the proximal thoracic spine.
  • 2.
    6 cases (16%) are due to over-smoothing of the upper thoracic curve under poor visibility or low image contrast.
  • 3.
    2 cases (6%) result from severe metal artifacts near the upper spine.
78% failure cases confirm that the main reason is the mislocalization of cervical landmarks, which extends the fitted curve cranially and leads to overestimation. To further validate this, we compute the average Cobb angle deviation when cervical landmarks are present in the predicted PT region: +14.3°, significantly larger than +3.1° when only thoracic landmarks (T1–T5) are detected.
Besides, it can be seen from Table 3 that segmentation-based method Residual U-Net achieves the best PT Cobb angle estimation result on Spinal2023, and landmark-based method Linformer achieves the best PT Cobb angle estimation result on Spinal-AI2024. We notice that Spinal2023 has limited images while Spinal-AI2024 has a larger scale. In this case, landmark-based methods suffer from insufficient training data on Spinal2023 but can be trained well on Spinal-AI2024, while segmentation-based methods have larger advantages on Spinal2023 with smaller scale. In contrast, our CurvNet consistently works well across small and large datasets.
@@ 4.3.1 Spinal Contour Detection Network
Our spinal contour detection network consists of latent contour coefficient regression branch and anchor box classification branch, with each branch containing sparse and dense assignments. Table 4 presents the results of different variants on Spinal2023.
@@
Table 5: Performance of our spinal contour detection network at different input image resolutions on AASCE2019.
Resolution SMAPE (%) FPS
$512\times 512$ 5.05 16.7
Native Solutions (Averagely $2299.5\times 894.3$) 5.31 5.6
Robustness on Input Image Resolutions. In our method, we resize input images to be $512\times 512$ to feed into the spinal contour detection network. To validate the robustness of our method on resolutions of input images, we also conduct an experiment by directly feeding the native images, as presented in Table 5. It can be observed that our spinal contour detection network still achieves comparable 5.31% SMAPE, demonstrating good robustness. Since the average resolution of native images from AASCE2019 is $2299.5\times 894.3$ and is significantly larger than $512\times 512$, the average inference speed is degraded to be 5.6 FPS. Therefore, we preserve the input $512\times 512$ image resize setting.
@@ 4.3.2 Data Engine
Table 6: Maximum Cobb angle estimation results of our spinal contour detection network using different hyperparameters on AASCE2019.
$\tau_{CS}=0.6$, $\lambda_{ss}/\lambda_{ps}=0.2/0.8$ $\tau_{c}$ 0.25 0.30 0.35
SMAPE (%) 5.42 5.05 5.38
$\tau_{CS}=0.6$, $\tau_{c}=0.3$ $\lambda_{ss}/\lambda_{ps}$ 0.1/0.9 0.2/0.8 0.5/0.5
SMAPE (%) 5.12 5.05 5.18
We also evaluate the sensitivity of key hyperparameters: the contour confidence threshold $\tau_{c}$, the consistency filtering threshold $\tau_{CS}$, and the loss weight ratio $\lambda_{ss}/\lambda_{ps}$, as shown in Table 6. We set $\tau_{CS}=0.6$ based on privacy-preserving criteria [13], where re-identification risk exceeds 80% when $\tau_{CS}>0.6$. It can be observed that SMAPE remains below 6% for $\tau_{c}\in[0.25,0.35]$, with optimal performance at $\tau_{c}=0.3$. For loss balancing, we use $\lambda_{ss}:\lambda_{ps}=0.2:0.8$ to prioritize pixel similarity related to $\mathcal{L}_{ps}$, which better preserves patient-specific anatomy, and achieves the lowest SMAPE.
@@
Table 8: Results using different selections on Spinal-AI2024 from the first to the sixth iterations. “No”, “Indep.”, and “Cumul.” denote no filtering, and updating $V_{selected}$ in independent and cumulative ways across iterations, respectively.
@@
Different Sample Selection Strategies. We conduct three experiments using different sample selection strategies, as shown in Table 8. It can be observed that the AP results of Cumul. stabilize after the fifth iteration, validating our early stopping strategy. If a sample is filtered and will not be selected in the next iterations, i.e. cumulative way, the overall highest quality improvement can be achieved. If $V_{selected}$ is updated independently across iterations, the quality improvement becomes worse but is still better than no filtering. This demonstrates the cumulative effectiveness of combined denoising and privacy review processes.
@@
Discussion on Spinal Image Generation Model. In our method, we adopt latent diffusion model over GAN due to the superior mode coverage and training stability. Additionally, diffusion model offers better privacy preservation by reducing memorization risk. To validate our image generation model, we train a StyleGAN2-ADA [40] model on Spinal2023 for comparison, and compute the Fréchet inception distance (FID) score between real Spinal2023 images and generated images using a pre-trained ResNet-50. The FID score of StyleGAN2-ADA is 41.3 with severe performance dropping in extreme cases like Cobb angle larger than 45°, which is significantly worse than 28.6 FID score of our diffusion based method.
@@ 5.2.3 Generated Spinal-AI2024
To assess the perceptual realism of the generated Spinal-AI2024 dataset, we compute the FID score between real images from Spinal2023 and generated images from Spinal-AI2024 using a pre-trained ResNet-50, resulting in a score of 28.6, indicating good visual fidelity. Moreover, three experts independently rate 100 randomly selected generated X-ray images on a five-point Likert scale:
  • 1.
    1: Clearly synthetic / unrealistic
  • 2.
    2: Mostly synthetic
  • 3.
    3: Ambiguous
  • 4.
    4: Mostly realistic
  • 5.
    5: Indistinguishable from real
The distribution of scores is summarized in Table 9. The average score is 4.3, with 82% of images scored $\geq 4$, confirming strong perceptual realism.
Table 9: Results of expert evaluation, in which three experts score 100 randomly selected generated images.
   Score    Description    Percentage
   1–2    Synthetic    3%
   3    Ambiguous    15%
   4    Mostly realistic    52%
   5    Indistinguishable    30%
Table 10: Image Similarity Review Checklist (excerpt). “New Image” denotes images from the generated dataset Spinal-AI2024, “Topkk-Image” denotes the image in the $k$-th place of descending similarity in the private dataset Spinal2023, and “ACS” denotes the average of Top1-CS, Top2-CS, and Top3-CS.
Figure 6: Illustration of highest similar images.
Besides, for each Spinal-AI2024 image, a top-3 comprehensive similarity review is presented in Table 10, in which the most similar images are shown in Fig. 6. The results show that our generated dataset has distorted the original annotations and produced spinal curvature variations, which can prevent excessive patient information leakage.
@@ 6 Conclusion
We have proposed a novel curvature angle estimation framework with latent contour representation based contour detection and iterative data engine based image self-generation. Latent contour coefficient regression and anchor box classification with sparse and dense assignments are combined to detect irregular contours of spinal segments. Besides, we have proposed an iterative data engine with semi-supervised labeling and privacy review, which culminates in the open-source largest scoliosis X-ray dataset Spinal-AI2024. Extensive experiments have demonstrated the effectiveness of our method, achieving state-of-the-art performance on public, private, and generated datasets for both maximum Cobb angle and three regional Cobb angles. In future work, we will improve PT Cobb angle estimation by introducing anatomical constraints or region-specific landmark filtering.
CRediT authorship contribution statement
Zhiwen Shao: Writing – review & editing, Methodology, Supervision, Funding acquisition, Conceptualization. Yichen Yuan: Writing - original draft, Visualization, Investigation, Software. Lizhuang Ma: Formal analysis, Validation, Funding acquisition. Xiaojia Zhu: Writing – review & editing, Data curation, Resources, Project administration.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62472424, 62472282, and 72192821, in part by the China Postdoctoral Science Foundation under Grant 2023M732223, and in part by the Hong Kong Scholars Program under Grant XJ2023037/HKSP23EG01.
Data availability
The used AASCE2019 and Spinal-AI2024 can be downloaded at https://aasce19.grand-challenge.org and https://github.com/Ernestchenchen/Spinal-AI2024, respectively. The used Spinal2023 is private due to privacy protection. Our code can be downloaded at https://github.com/Ernestchenchen/CurvNet.